OpenStudy (anonymous):

Which is the interquartile range of the data set: 6, 8, 10, 12, 14?

OpenStudy (anonymous):

I got 8=Q1, 12=Q2 AND THEN i ADDED 12+8=20/2=10

Directrix (directrix):

The IQR is the value of the third quartile minus the first quartile: Q3 - Q1. I got Q3 - Q1 = 13 - 7 = 6

OpenStudy (anonymous):

what did you get for Q1 and Q2? thanks for showing me.

Directrix (directrix):

Q1 = 7 and Q3 = 13 @fdrsgdfgdfg